# xDeepServe: Model-as-a-Service on Huawei CloudMatrix384

### v5

@@ 5.1 Disaggregated Prefill and Decode
    Upon completion of prefill, the DP master registers a PD-transfer task with RTC-DistFlow hu2025deepflow . This task contains only metadata and the addresses of the relevant KV cache blocks; actual transfer is deferred until triggered by the decode phase.
@@
    On the decode side, the DP checks for available KV block slots. If capacity is insufficient, the RECV is deferred, applying backpressure to upstream components. If capacity is sufficient, an asynchronous RECV is submitted to DistFlow. We use the point-to-point send/recv APIs described in §3.1.
@@
    DistFlow orchestrates the actual KV data transfer, handling all low-level concerns such as SEND/RECV handshakes, ordering guarantees, and TP rank synchronization. Since KV blocks are not self-describing, DistFlow ensures correct semantic pairing between prefill and decode DPs. Each prefill–decode TE pair operates with an isolated DistFlow instance to limit failure domains. However, multiple DistFlow instances can share XCCL buffers to optimize NPU’s on-chip memory usage.
@@
Heterogeneous Prefill and Decode Deployment. To maximize cost-efficiency, we run prefill and decode on heterogeneous NPUs. Prefill runs on both scale-out Ascend 910 (the second generation) and scale-up CloudMatrix384 NPUs, since it is compute-bound and largely indifferent to interconnect bandwidth. We place decode exclusively on CloudMatrix384, where high-speed interconnect is key to meeting the 35 ms TPOT SLA by accelerating MoE dispatch and combine. When prefill runs on Ascend 910 (the second generation) and decode on CloudMatrix384, we transfer KV cache over RoCE or VPC interconnects, as shown in Figure 2. FlowServe selects the appropriate DistFlow hu2025deepflow backend based on the network fabric. For MLA models like DeepSeek and Kimi K2, both interconnects satisfy TTFT and TPOT SLAs.
5.2 Disaggregated MoE and Attention
Figure 17: The Archiecture of Disaggregated MoE and Attention over CloudMatrix384. Our deployment spans a full SuperPod with 768 NPU dies: 288 run EP288 (256 routed experts and 32 shared experts), and 480 handle MLA computation. We deploy three DP domains, each with 160 DP groups and TP=1. Attention DP groups execute full MLA computation, including MLAProlog, Attention, gating, output projection, and A2E/E2A kernels. MoE NPUs run only A2E, MoE computation, and E2A.
Figure 18: The Execution Pipeline of Disaggregated MoE and Attention. Note that each DP domain has 160 DP groups with TP set to 1. XCCL A2E and E2A uses two-stage routing. All 288 MoE NPU run that same persistent kernels with three concurrent streams.
The idea of disaggregating attention from feedforward networks (FFN) and executing them on separate resources was pioneered by works such as FastDecode he2024fastdecode , Lamina chen2024efficient-ma-mingxing , and InstAttention pan2025instattention , and later extended to MoE models by MegaScale-Infer megascale-infer . However, prior work primarily explores small models at limited scale, leaving it unclear whether the same principles hold for large-scale MoE models like DeepSeek.
To bridge this gap, we build the first large-scale disaggregated MoE-Attention system for MoE models over CloudMatrix384. Specifically, we run DeepSeek-V3/R1 with MoE-Attention disaggregated at the decode stage. Our deployment spans a full CloudMatrix384 with 768 NPU dies: 288 dies run EP288 (256 routed experts and 32 shared experts), and 480 dies runs MLA. We use DeepSeek as an example, the same design works for Kimi K2 kimi-k2 and Qwen yang2025qwen3technicalreport .
@@ 5.2 Disaggregated MoE and Attention
To solve these challenges, we introduce three techniques illustrated in Figures 17 and 18.
@@ 6.2 Failure Recovery
Stage 3: Fine-Grained Error Handling. Many transient failures don’t require engine restarts. We focus on two frequent cases: network instability and on-chip memory faults. Network glitches (e.g., from switch flapping or BGP convergence) may temporarily break communication. To tolerate this, FlowServe adopts a token recomputation strategy. When certain error codes are detected, all DP groups coordinate a rollback to the previous iteration. A dedicated thread broadcasts the rollback signal to ensure all groups—including those in busy-wait—are notified. The iteration is then re-executed, avoiding full restarts. On-chip memory faults are another common issue. When triggered, we work with the CANN runtime to remap virtual memory and mask the faulty region. While this may cause some KV cache loss and individual request failures, the system remains online and continues serving unaffected requests.
@@ 7.1 Decode Performance
We evaluate the maximum decoding performance of running DeepSeek liu2024deepseek-v3 on CloudMatrix384.
Setup. We evaluate FlowServe on a single CloudMatrix384 deployment comprising 18 Ascend servers, totaling 288 NPU dies, using a colocated prefill-and-decode setup. The system is configured with DP288 and EP288: 256 dies host one routed expert and one redundant expert each, while the remaining 32 dies host one shared expert and one redundant expert each. We enable MTP with one MTP layer. Each DP/die handles a local batch size of 60, yielding a global batch size of $60\times 288=17{,}280$. FlowServe is configured to complete prefill for all requests before entering the decode phase simultaneously. The workload uses fixed 2K-token prompts manually constructed from ShareGPT sharegpt , followed by 2K-token outputs. We enforce fixed-length outputs by enabling ignore-eos. Both the MTP module and the main model use greedy sampling during inference.
Performance. We achieve an average TPOT (time per output token) of 50 ms. Each decode iteration—which includes MTP forward, sampling, main model execution, and final sampling (see Figure 13)—takes approximately 93 ms. The scheduling gap between iterations is around 2 ms. With an MTP acceptance rate of about 90% across all DPs, the effective TPOT is computed as $\frac{93+2}{1.9}\approx 50$ ms. This translates to a per-chip throughput of $2\times 60\times\frac{1000}{50}=2400$ tokens/second (two dies per chip, each with batch size 60). For the full DP288 system, the total throughput reaches 345K tokens/second. Figure 19 shows kernel-level latency breakdown for one decode iteration (at approximately 3K sequence length). The attention kernel—scaling with both sequence and batch size—accounts for 21.8% of latency and will grow with longer sequences. Thanks to CloudMatrix384’s high-bandwidth UB fabric, dispatch and combine are efficient, contributing about 36% of the total latency. However, these global synchronization kernels exhibit significant variance: their maximum latency can be up to 10× their minimum. This is because dispatch absorbs variance from MLA compute across all DPs, while combine absorbs imbalance from MoE compute across experts (see Figure 10 for execution timeline).
Disaggregated MoE and Attention. We evaluate DeepSeek-R1’s decoding performance using disaggregated MoE-Attention on a single CloudMatrix384 comprising 768 NPU dies. Prefill is executed separately on other NPU servers. The decode TE is configured as follows: 480 NPU dies are organized into three DP domains, each with 160 DP groups. The remaining 288 NPU dies run EP288. Each DP/die processes a local batch size of 96, resulting in a global batch size of $96\times 3\times 160=46{,}080$. We use a fixed 2K+2K workload with ignore-eos enabled and greedy sampling. Under this setup, we achieve a per-chip throughput of 2400 tokens/second with a 50 ms TPOT. Specifically, MLAProlog, MLA, Gating, and the first stage of A2E each take approximately 700 ns per layer. TPOT includes both MTP and main model forwards; the second microbatch of the final layer cannot be overlapped. Accounting for 2 ms scheduler overhead, 5 ms for the MTP layer, and layer-wise compute $(0.7\times 2\times 61)$ ms, plus latencies for A2E (0.17 ms), MoE (0.12 ms), and E2A (0.19 ms), the total forward time is approximately 93 ms. Given an MTP acceptance rate of 90% across all DPs, the effective TPOT is $\frac{93}{1.9}\approx 49$ ms.
@@ 7.2 Production Workload
We now report the performance of a representative production setup and workload. The deployment consists of 16 Ascend servers, organized into four prefill TEs and one decode TE. Each prefill TE spans two servers and is configured with DP8 and EP32. The decode TE spans eight servers with DP128 and EP128. The workload features input lengths ranging from 0 to 64K tokens, with an average input length of 13K and an average output length of 2.1K tokens. Under this setup, we achieve a TTFT of 900 ms and an average TPOT of 34.8 ms.
@@ 9 Contributors
Ao Xiao, Bangzheng He, Baoquan Zhang, Baoxing Huai, Bingji Wang, Bo Wang, Bo Xu, Boyi Hou, Chan Yang, Changhong Liu, Cheng Cui, Chenyu Zhu, Cong Feng, Daohui Wang, Dayun Lin, Duo Zhao, Fengshao Zou, Fu Wang, Gangqiang Zhang, Gengyuan Dan, Guanjie Chen, Guodong Guan, Guodong Yang, Haifeng Li, Haipei Zhu, Haley Li, Hao Feng, Hao Huang, Hao Xu, Hengrui Ma, Hengtao Fan, Hui Liu, Jia Li, Jiang Liu, Jiang Xu, Jie Meng, Jinhan Xin, Junhao Hu, Juwei Chen, Lan Yu, Lanxin Miao, Liang Liu, Linan Jing, Lu Zhou, Meina Han, Mingkun Deng, Mingyu Deng, Naitian Deng, Nizhong Lin, Peihan Zhao, Peng Pan, Pengfei Shen, Ping Li, Qi Zhang, Qian Wang, Qin ZhC Qingrong Xia, Qingyi Zhang, Qunchao Fu, Ren Guo, Ruimin Gao, Shaochun Li, Sheng Long, Shentian Li, Shining Wan, Shuai Shen, Shuangfu Zeng, Shuming Jing, Siqi Yang, Song Zhang, Tao Xu, Tianlin Du, Ting Chen, Wanxu Wu, Wei Jiang, Weinan Tong, Weiwei Chen, Wen Peng, Wenli Zhou, Wenquan Yang, Wenxin Liang, Xiang Liu, Xiaoli Zhou, Xin Jin, Xinyu Duan, Xu Li, Xu Zhang, Xusheng Chen, Yalong Shan, Yang Gan, Yao Lu, Yi Deng, Yi Zheng, Ying Xiong, Yingfei Zheng, Yiyun Zheng, Yizhou Shan, Yong Gao, Yong Zhang, Yongqiang Yang, Yuanjin Gong, Yue Yu, Yuetao Chen, Yukun Zhu, Yulong He, Yusu Zhao, Yuyan Wu, Zenan Zhang, Zhaojin Zhuo, Zhaoyang Ji, Zhefeng Wang, Zheng Wang, Zhenan Fan, Zhenhua Yang, Zhenli Sheng, Zhibin Yu, Zhigang Ji, Zhihao Ren, Zhipeng Bian, Zhixia Liu, Zhiyu Dong, Zhonghua Li, Zhou Yu, Zhuoming Shen, Zhuwei Peng, Zi Ye, Zihao Xiang, Zimin Fu, Zixuan Zhang.

### v6

@@ 5.1 Disaggregated Prefill-Decode
    Upon completion of prefill, the DP master registers a PD-transfer task with RTC-DistFlow [hu2025deepflow]. This task contains only metadata and the addresses of the relevant KV cache blocks; actual transfer is deferred until triggered by the decode phase.
@@
    On the decode side, the DP checks for available KV block slots. If capacity is insufficient, the RECV is deferred, applying backpressure to upstream components. If capacity is sufficient, an asynchronous RECV is submitted to DistFlow. We use the point-to-point send/receive APIs described in §3.1.
@@
    DistFlow orchestrates the actual KV data transfer, handling all low-level concerns such as SEND/RECV handshakes, ordering guarantees, and TP rank synchronization. Since KV blocks are not self-describing, DistFlow ensures correct semantic pairing between prefill and decode DPs. Each prefill-decode TE pair operates with an isolated DistFlow instance to limit failure domains. However, multiple DistFlow instances can share XCCL buffers to optimize NPU’s on-chip memory usage.
@@
Heterogeneous Prefill-Decode Deployment. To maximize cost-efficiency, we run prefill and decode on heterogeneous NPUs. Prefill runs on both scale-out Ascend 910B and scale-up Ascend 910C, since it is compute-bound and largely indifferent to interconnect bandwidth. We place decode exclusively on Ascend 910C CloudMatrix384, where high-speed interconnect is key to meeting the 35 ms TPOT SLA by accelerating MoE dispatch and combine. When prefill runs on 910B and decode on 910C, we transfer KV cache over RoCE or VPC interconnects, as shown in Figure 2. FlowServe selects the appropriate DistFlow [hu2025deepflow] backend based on the network fabric. For MLA models like DeepSeek and Kimi K2, both interconnects satisfy TTFT and TPOT SLAs.
5.2 Disaggregated MoE-Attention
Figure 18: The Architecture of Disaggregated MoE-Attention over CloudMatrix384. Our deployment spans a full SuperPod with 768 NPU dies: 288 run EP288 (256 routed experts and 32 shared experts), and 480 handle MLA computation. We deploy three DP domains, each with 160 DP groups and TP=1. Attention DP groups execute full MLA computation, including MLAProlog, Attention, gating, output projection, and A2E/E2A kernels. MoE NPUs run only A2E, MoE computation, and E2A.
Figure 19: The Execution Pipeline of Disaggregated MoE-Attention. Note that each DP domain has 160 DP groups with TP set to 1. XCCL A2E and E2A use two-stage routing. All 288 MoE NPUs run the same persistent kernels with three concurrent streams.
The idea of disaggregating attention from feedforward networks (FFN) and executing them on separate resources was pioneered by works such as FastDecode [he2024fastdecode], Lamina [chen2024efficient-ma-mingxing], and InstAttention [pan2025instattention], and later extended to MoE models by MegaScale-Infer [megascale-infer]. However, prior work primarily explores small models at limited scale, leaving it unclear whether the same principles hold for large-scale MoE models like DeepSeek.
To bridge this gap, we build a large-scale disaggregated MoE-Attention system for MoE models over CloudMatrix384. Specifically, we run DeepSeek-V3/R1 with MoE-Attention disaggregated at the decode stage. Our deployment spans a full CloudMatrix384 with 768 NPU dies: 288 dies run EP288 (256 routed experts and 32 shared experts), and 480 dies run MLA. We use DeepSeek as an example, the same design works for Kimi K2 [kimi-k2] and Qwen [yang2025qwen3technicalreport].
@@ 5.2 Disaggregated MoE-Attention
To solve these challenges, we introduce three techniques illustrated in Figures 18 and 19.
@@ 6.2 Failure Recovery
Stage 3: Fine-Grained Error Handling. Many transient failures don’t require engine restarts. We focus on two frequent cases: network instability and on-chip memory faults. Network glitches (e.g., from switch flapping or BGP convergence) may temporarily break communication. To tolerate this, FlowServe adopts a token recomputation strategy. When certain error codes are detected, all DP groups coordinate a rollback to the previous iteration. A dedicated thread broadcasts the rollback signal to ensure all DP groups—including those in busy-wait—are notified. The iteration is then re-executed, avoiding full restarts. On-chip memory faults are another common issue. When triggered, we work with the CANN runtime to remap virtual memory and mask the faulty region. While this may cause some KV cache loss and individual request failures, the system remains online and continues serving unaffected requests.
@@ 7.1 Decode Performance
We evaluate the maximum decoding performance of running DeepSeek [liu2024deepseek-v3] on CloudMatrix384.
Setup. We evaluate FlowServe on a single CloudMatrix384 deployment comprising 18 Ascend 910C servers, totaling 288 NPU dies, using a colocated prefill-and-decode setup. The system is configured with DP288 and EP288: 256 dies host one routed expert and one redundant expert each, while the remaining 32 dies host one shared expert and one redundant expert each. We enable MTP with one MTP layer. Each DP/die handles a local batch size of 60, yielding a global batch size of $60\times 288=17{,}280$. FlowServe is configured to complete prefill for all requests before entering the decode phase simultaneously. The workload uses fixed 2K-token prompts manually constructed from ShareGPT [sharegpt], followed by 2K-token outputs. We enforce fixed-length outputs by enabling ignore-eos. Both the MTP module and the main model use greedy sampling during inference.
[…]
Disaggregated MoE-Attention. We evaluate DeepSeek-R1’s decoding performance using disaggregated MoE-Attention on a single CloudMatrix384 comprising 768 NPU dies. Prefill is executed separately on other NPU servers. The decode TE is configured as follows: 480 NPU dies are organized into three DP domains, each with 160 DP groups. The remaining 288 NPU dies run EP288. Each DP/die processes a local batch size of 96, resulting in a global batch size of $96\times 3\times 160=46{,}080$. We use a fixed 2K+2K workload with ignore-eos enabled and greedy sampling. Under this setup, we achieve a per-chip throughput of 2400 tokens/second with a 50 ms TPOT. Specifically, MLAProlog, MLA, Gating, and the first stage of A2E each take approximately 700 ns per layer. TPOT includes both MTP and main model forwards; the second microbatch of the final layer cannot be overlapped. Accounting for 2 ms scheduler overhead, 5 ms for the MTP layer, and layer-wise compute $(0.7\times 2\times 61)$ ms, plus latencies for A2E (0.17 ms), MoE (0.12 ms), and E2A (0.19 ms), the total forward time is approximately 93 ms. Given an MTP acceptance rate of 90% across all DPs, the effective TPOT is $\frac{93}{1.9}\approx 49$ ms.
@@ 7.2 Production Workload
We now report the performance of a representative production setup and workload. The deployment consists of 16 Ascend 910C servers, organized into four prefill TEs and one decode TE. Each prefill TE spans two servers and is configured with DP8 and EP32. The decode TE spans eight servers with DP128 and EP128. The workload features input lengths ranging from 0 to 64K tokens, with an average input length of 13K and an average output length of 2.1K tokens. Under this setup, we achieve a TTFT of 900 ms and an average TPOT of 34.8 ms.
@@ 9 Contributors
Ao Xiao, Bangzheng He, Baoquan Zhang, Baoxing Huai, Bingji Wang, Bo Wang, Bo Xu, Boyi Hou, Chan Yang, Changhong Liu, Cheng Cui, Chenyu Zhu, Cong Feng, Daohui Wang, Dayun Lin, Duo Zhao, Fengshao Zou, Fu Wang, Gangqiang Zhang, Gengyuan Dan, Guanjie Chen, Guodong Guan, Guodong Yang, Haifeng Li, Haipei Zhu, Hao Feng, Hao Huang, Hao Xu, Hengrui Ma, Hengtao Fan, Hui Liu, Jia Li, Jiang Liu, Jiang Xu, Jie Meng, Jinhan Xin, Junhao Hu, Juwei Chen, Lan Yu, Lanxin Miao, Liang Liu, Linan Jing, Lu Zhou, Meina Han, Mingkun Deng, Mingyu Deng, Naitian Deng, Nizhong Lin, Peihan Zhao, Peng Pan, Pengfei Shen, Ping Li, Qi Zhang, Qin Zhang, Qingrong Xia, Qingyi Zhang, Qunchao Fu, Ren Guo, Ruimin Gao, Shaochun Li, Sheng Long, Shentian Li, Shining Wan, Shuai Shen, Shuangfu Zeng, Shuming Jing, Siqi Yang, Song Zhang, Tao Xu, Tianlin Du, Ting Chen, Wanxu Wu, Wei Jiang, Weinan Tong, Weiwei Chen, Wen Peng, Wenli Zhou, Wenquan Yang, Wenxin Liang, Xiang Liu, Xiaoli Zhou, Xin Jin, Xinyu Duan, Xu Li, Xu Zhang, Xusheng Chen, Yalong Shan, Yang Gan, Yao Lu, Yi Deng, Yi Zheng, Yingfei Zheng, Yiyun Zheng, Yizhou Shan, Yong Gao, Yongqiang Yang, Yuanjin Gong, Yue Yu, Yuetao Chen, Yukun Zhu, Yulong He, Yusu Zhao, Yuyan Wu, Zenan Zhang, Zhaojin Zhuo, Zhaoyang Ji, Zhefeng Wang, Zheng Wang, Zhenhua Yang, Zhenli Sheng, Zhibin Yu, Zhigang Ji, Zhihao Ren, Zhipeng Bian, Zhixia Liu, Zhiyu Dong, Zhonghua Li, Zhou Yu, Zhuoming Shen, Zhuwei Peng, Zi Ye, Zihao Xiang, Zimin Fu, Zixuan Zhang.